the complete combustion of octane , C8H18, a component og gasoline, proceed as 2 C8H18 +25 O2 = 16 CO2 + 18H2O.
a) how many moles of CO2 are produced when 1.50 mol octance reacted?
b) how many grams of water produced in this reaction?
c)how many moles of oxygen required to form 90.0 g water?
Answer #1

a) 2 moles of octane gives 16 moles of CO2.

1 mole of octane will give 8 moles of CO2

1.5 moles of octane will give = 1.5 × 8 moles of CO2.

Moles of CO2 formed = 12 moles CO2

b.) Number of moles = Mass / molar mass

Molar mass of H2O = 18 g/mol

In this reaction 18 moles of water are formed.

Mass of water formed = 18 × 18

= 324 grams of water.

C.) 25 moles of oxygen reacts to give 18 moles of water.

1 mole of water will be formed by 25/18 moles of oxygen

Number of moles = Mass / molar mass

Moles of water = 90 / 18

Moles of water = 5 moles.

Moles of oxygen required =( 25/18) × 5

Moles of oxygen required = 6.94 moles